# Construction of 2-D Z-Complementary Array Code Sets with Flexible Lengths for Different System Requirements

### v3

@@ Appendix A Proof of Lemma 5
$\begin{split}&C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)(0)\\ &=\sum_{v^{\prime}_{p_{1}}=0}^{p_{1}-1}\sum_{v^{\prime}_{p_{2}}=0}^{p_{2}-1}\dots\sum_{v^{\prime}_{p_{k}}=0}^{p_{k}-1}\omega_{q}^{\sum_{\alpha=1}^{k}\frac{q}{p_{\alpha}}v^{\prime}_{p_{\alpha}}(s^{1}_{\alpha}-s^{2}_{\alpha})}\\ &=\sum_{v^{\prime}_{p_{1}}=0}^{p_{1}-1}\omega_{p_{1}}^{v^{\prime}_{p_{1}}(s^{1}_{1}-s^{2}_{1})}\times\dots\times\sum_{v^{\prime}_{p_{k}}=0}^{p_{k}-1}\omega_{p_{k}}^{v^{\prime}_{p_{k}}(s^{1}_{k}-s^{2}_{k})}.\end{split}$ (51)
@@
From (51), (52) for $\mathbf{s}_{1}=\mathbf{s}_{2}$ we have
@@
Now we come to the main proof of Lemma 5. Note that
$\psi(a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}})=\psi(T_{\mathbf{s}})\otimes\psi(R_{\mathbf{t}}^{\bm{\gamma}}),$ (55)
where $\otimes$ denotes the Kronecker product. Then we have
$\begin{split}&\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}})\right)(\tau)\\ &=\bigg{[}C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)\left(\left\lfloor\frac{\tau}{N}\right\rfloor\right)\\ &\quad\times\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\mod N\right)\bigg{]}\\ &\quad+\bigg{[}\Delta_{N}C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)\left(\left\lfloor\frac{\tau}{N}\right\rfloor+1\right)\\ &\quad\times\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\mod N-N\right)\bigg{]},\end{split}$ (56)
with $N=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$, because for two pairs of sequences $(\mathbf{a}_{1},\mathbf{a}_{2})$, $(\mathbf{b}_{1},\mathbf{b}_{2})$ and $\tau\geq 0$, we have
$\begin{split}&C(\mathbf{a}_{1}\otimes\mathbf{b}_{1},\mathbf{a}_{2}\otimes\mathbf{b}_{2})(\tau)\\ &=C(\mathbf{a}_{1},\mathbf{a}_{2})\left(\left\lfloor\frac{\tau}{N}\right\rfloor\right)C(\mathbf{b}_{1},\mathbf{b}_{2})(\tau\!\!\!\!\mod N)\\ &\quad+\!\Delta_{N}C(\mathbf{a}_{1},\mathbf{a}_{2})\!\left(\left\lfloor\frac{\tau}{N}\right\rfloor\!\!+\!\!1\right)\!C(\mathbf{b}_{1},\mathbf{b}_{2})(\tau\!\!\!\!\!\mod N\!\!-\!\!N),\end{split}$ (57)
where each $\mathbf{a}_{i}$ has length $M$ and each $\mathbf{b}_{i}$ has length $N$, and
@@
We need to show that
$\begin{split}&\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}})\right)(\tau)\\ &=\begin{cases}\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}+1},&\left(\mathbf{s}_{1},\mathbf{t}_{1}\right)=\left(\mathbf{s}_{2},\mathbf{t}_{2}\right),\tau=0;\\ 0,&\left(\mathbf{s}_{1},\mathbf{t}_{1}\right)=\left(\mathbf{s}_{2},\mathbf{t}_{2}\right),\\ &0<\tau<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1};\\ 0,&\mathbf{s}_{1}\neq\mathbf{s}_{2},\mathbf{t}_{1}=\mathbf{t}_{2},\tau<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}.\\ \end{cases}\end{split}$ (59)
@@ 1. Case I:
    $\begin{split}&\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}})\right)(\tau)\\ &=\begin{cases}\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}+1},&\mathbf{s}_{1}=\mathbf{s}_{2},\mathbf{t}_{1}=\mathbf{t}_{2};\\ 0,&\mathbf{s}_{1}\neq\mathbf{s}_{2},\mathbf{t}_{1}=\mathbf{t}_{2};\\ 0,&\mathbf{t}_{1}\neq\mathbf{t}_{2}.\end{cases}\end{split}$ (60)
@@ 2. Case II:
    Let $0<\tau<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$. In this case $\left\lfloor\frac{\tau}{N}\right\rfloor=0$, $0<\tau\mod N<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$, $-\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau\mod N-N<0$ and $\Delta_{N}=1$. Using Lemma 9, we have
@@ 2. Sub-case 2:
      Let $\beta\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau<(\beta+1)\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$ for some $\beta\in\{1,2,\dots,\prod_{\alpha=1}^{k}p_{\alpha}-1\}$. Then $\left\lfloor\frac{\tau}{N}\right\rfloor=\beta$, $0<\tau\mod N<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$, and $-\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau\mod N-N<0$ with $\Delta_{N}=1$. Then using Lemma 9, we have
@@
Hence, the lemma is proved.
Appendix B Proof of Theorem 1
@@
$\begin{split}&\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})\\ &=\begin{cases}L_{1}L_{2}\prod_{\alpha=1}^{k}p_{\alpha},&\bm{\zeta}_{1}=\bm{\zeta}_{2},(\tau_{1},\tau_{2})=(0,0);\\ 0,&\bm{\zeta}_{1}=\bm{\zeta}_{2},(\tau_{1},\tau_{2})\neq(0,0);\\ &\left|\tau_{1}\right|<2^{m+1},\left|\tau_{2}\right|<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1};\\ 0,&\bm{\zeta}_{1}\neq\bm{\zeta}_{2},\left|\tau_{1}\right|<2^{m+1},\\ &\left|\tau_{2}\right|<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1},\\ \end{cases}\end{split}$ (65)
@@
$\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})=2^{m+1}p\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}+1},\end{split}$ (66)
@@
$\begin{split}&C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})\\ &=\sum_{i=0}^{L_{1}-\tau_{1}-1}\sum_{j=0}^{L_{2}-\tau_{2}-1}\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}}_{i,j}{\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}_{i^{\prime},j^{\prime}}}^{*}\\ &=\sum_{i=0}^{L_{1}-\tau_{1}-1}\sum_{j=0}^{L_{2}-\tau_{2}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i^{\prime}}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j^{\prime}})},\end{split}$ (67)
where $i^{\prime}=i+\tau_{1}$, $j^{\prime}=j+\tau_{2}$. Let $1+\sum_{\alpha=1}^{k}m_{\alpha}=\epsilon$ and $m+1=\vartheta$. We use the following notations.
  1. (1)
@@
$\begin{split}&\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j}^{\prime})}\\ &=\omega_{q}^{(1-j_{\epsilon})\left\{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})+a_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+a_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})\right\}}\\ &\quad\times\omega_{q}^{j_{\epsilon}\left\{a_{\mathbf{s}_{2}^{1},\mathbf{t}_{2}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})+b_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+b_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})\right\}}\\ &\quad\times\omega_{q}^{-(1-j^{\prime}_{\epsilon})\left\{a_{\mathbf{s}_{1}^{2},\mathbf{t}_{1}^{2}}^{\bm{\gamma}}(\vec{j}^{\prime}_{\xi})+a_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+a_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right\}}\\ &\quad\times\omega_{q}^{-j^{\prime}_{\epsilon}\left\{a_{\mathbf{s}_{2}^{2},\mathbf{t}_{2}^{2}}^{\bm{\gamma}}(\vec{j}^{\prime}_{\xi})+b_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+b_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right\}},\end{split}$ (68)

### v4

@@ 3.1 IGC Code Set
$\begin{split}C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)(0)=&\sum_{v^{\prime}_{p_{1}}=0}^{p_{1}-1}\sum_{v^{\prime}_{p_{2}}=0}^{p_{2}-1}\dots\sum_{v^{\prime}_{p_{k}}=0}^{p_{k}-1}\omega_{q}^{\sum_{\alpha=1}^{k}\frac{q}{p_{\alpha}}v^{\prime}_{p_{\alpha}}(s^{1}_{\alpha}-s^{2}_{\alpha})}\\ =&\sum_{v^{\prime}_{p_{1}}=0}^{p_{1}-1}\omega_{p_{1}}^{v^{\prime}_{p_{1}}(s^{1}_{1}-s^{2}_{1})}\times\dots\times\sum_{v^{\prime}_{p_{k}}=0}^{p_{k}-1}\omega_{p_{k}}^{v^{\prime}_{p_{k}}(s^{1}_{k}-s^{2}_{k})}.\end{split}$ (33)
@@
From (33) and (34), for $\mathbf{s}_{1}=\mathbf{s}_{2}$ we have
@@
Now, we can prove Lemma 5. We consider the following observation before proving the lemma. Note that for two pairs of sequences $(\mathbf{a}_{1},\mathbf{a}_{2})$, $(\mathbf{b}_{1},\mathbf{b}_{2})$ and $\tau\geq 0$, we have
$\begin{split}C(\mathbf{a}_{1}\otimes\mathbf{b}_{1},\mathbf{a}_{2}\otimes\mathbf{b}_{2})(\tau)=&C(\mathbf{a}_{1},\mathbf{a}_{2})\left(\left\lfloor\frac{\tau}{N}\right\rfloor\right)C(\mathbf{b}_{1},\mathbf{b}_{2})(\tau\mod N)\\ &+\Delta_{N}C(\mathbf{a}_{1},\mathbf{a}_{2})\left(\left\lfloor\frac{\tau}{N}\right\rfloor+1\right)C(\mathbf{b}_{1},\mathbf{b}_{2})(\tau\mod N-N),\end{split}$ (37)
where $\otimes$ denotes the Kronecker product, each $\mathbf{a}_{i}$ has length $M$ and each $\mathbf{b}_{i}$ has length $N$, and
@@
Proof of Lemma 5.
We can write $\psi(a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}})$ as
$\psi(a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}})=\psi(T_{\mathbf{s}})\otimes\psi(R_{\mathbf{t}}^{\bm{\gamma}}).$ (39)
Then we have
$\begin{split}&\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}})\right)(\tau)\\ &=\bigg{[}C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)\left(\left\lfloor\frac{\tau}{N}\right\rfloor\right)\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\mod N\right)\bigg{]}\\ &\quad+\bigg{[}\Delta_{N}C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)\left(\left\lfloor\frac{\tau}{N}\right\rfloor+1\right)\times\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\!\!\!\!\!\!\!C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\mod N-N\right)\bigg{]},\end{split}$ (40)
where $N=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$. We need to show that
[…]
@@ 1. Case I:
    $\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}})\right)(\tau)=&\begin{cases}\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}+1},&\mathbf{s}_{1}=\mathbf{s}_{2},\mathbf{t}_{1}=\mathbf{t}_{2};\\ 0,&\mathbf{s}_{1}\neq\mathbf{s}_{2},\mathbf{t}_{1}=\mathbf{t}_{2};\\ 0,&\mathbf{t}_{1}\neq\mathbf{t}_{2}.\end{cases}\end{split}$ (42)
@@ 2. Case II:
    Let $0<\tau<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$. In this case $\left\lfloor\frac{\tau}{N}\right\rfloor=0$, $0<\tau\mod N<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$, $\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau\mod N-N<0$ and $\Delta_{N}=1$. Using Lemma 6, we have
@@ 2. Sub-case 2:
      Let $\beta\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau<(\beta+1)\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$ for some $\beta\in\{1,2,\dots,\prod_{\alpha=1}^{k}p_{\alpha}-1\}$. Then $\left\lfloor\frac{\tau}{N}\right\rfloor=\beta$, $0<\tau\mod N<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$, and $\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau\mod N-N<0$ with $\Delta_{N}=1$. Then using Lemma 6, we have
@@
Hence, the lemma is proved. ∎
We illustrate Lemma 5 by using the following example.
Example 2.
3.2 Main Construction
Now, we come to the main construction of 2-D ZCAC and 2-D ZCACS using EBFs obtained from the construction of the IGC code set. Before that, we state two more results which are used in the construction.
Lemma 8 ([paterson]).
where $\pi$ is a permutation of the set $\{1,2,\dots,m\}$ and $g_{\beta}\in\mathbb{Z}_{q}$. Then the complex-valued sequences $\mathbf{a}_{1}$ and $\mathbf{b}_{1}$, corresponding to the functions
$\begin{split}a_{1}=&f,\\ b_{1}=&f+\frac{q}{2}x_{\pi(1)},\end{split}$ (48)
respectively, forms a GCP of length $2^{m}$ for any $e\in\mathbb{Z}_{q}$.
Lemma 9 ([RathinaCCC]).
Let $f$, $a_{1}$ and $b_{1}$ be defined as in Lemma 8. Then the complex-valued sequences $\mathbf{a}_{2}$ and $\mathbf{b}_{2}$, corresponding to the functions
$\begin{split}a_{2}=&\bar{f}+\frac{q}{2}x_{\pi(1)},\\ b_{2}=&\bar{f},\end{split}$ (49)
respectively, forms a GCP of length $2^{m}$, where
For further development, we let $2\mid q$ and $p$ be a prime such that $p\mid q$. Also, let $a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}}$ be the same functions used in Lemma 5. For $\vec{d}^{\prime}=(d_{1},d_{2},\dots,d_{m},\tilde{d}_{p})\in\mathbb{Z}_{2}^{m}\times\mathbb{Z}_{p}$, $\mathbf{t}_{1}\neq\mathbf{t}_{2}$, and $\bm{\zeta}=(\mathbf{s}_{1},\mathbf{s}_{2},\mathbf{t}_{1},\mathbf{t}_{2})$, we define a set of functions
given by
$\begin{split}F_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}=&\bar{v^{\prime\prime}}\left\{a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}+a_{1}(\vec{d})W_{1}(\tilde{d}_{p})+a_{2}(\vec{d})W_{2}(\tilde{d}_{p})\right\}\\ &+v^{\prime\prime}\left\{a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}}+b_{1}(\vec{d})W_{1}(\tilde{d}_{p})+b_{2}(\vec{d})W_{2}(\tilde{d}_{p})\right\},\end{split}$ (52)
where
[…]
and
[…]
Note 4.
In the function $F_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}$, we could take the sub-function of the form (53) only after we have extended the co-domain of EBF beforehand from $\mathbb{Z}_{q}$ to $\mathbb{Q}_{q}$.
The only thing that now remains is to determine the values of the parameter $\bm{\zeta}$, which will produce the set size of 2-D ZCACS. Let us define the set $\Omega=\{(\mathbf{s}_{1},\mathbf{s}_{2},\mathbf{t}_{1},\mathbf{t}_{2}):\mathbf{t}_{1}\neq\mathbf{t}_{2}\}$. We take a proper subset $\Gamma\subset\Omega$ such that for any two $\bm{\zeta}_{1}=(\mathbf{s}_{1}^{1},\mathbf{s}_{2}^{1},\mathbf{t}_{1}^{1},\mathbf{t}_{2}^{1}),\bm{\zeta}_{2}=(\mathbf{s}_{1}^{2},\mathbf{s}_{2}^{2},\mathbf{t}_{1}^{2},\mathbf{t}_{2}^{2})\in\Gamma$, the terms $\mathbf{s}_{i}^{m},\mathbf{t}_{j}^{n}$ satisfy the following condition:
  1. 1.
    Either, $\mathbf{t}_{i}^{1}\neq\mathbf{t}_{j}^{2}$.
  2. 2.
    Or, $\mathbf{s}_{i}^{1}\neq\mathbf{s}_{j}^{2}$.
Following this, we consider only those functions $\mathcal{F}_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}$ for which $\bm{\zeta}\in\Gamma$. Let the $\mathbb{Q}_{q}$-valued 2-D arrays be given by
$\mathcal{K}^{\bm{\gamma},\bm{\zeta}}=\{F_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}:\vec{d}^{\prime}\in\mathbb{Z}_{2}^{m}\times\mathbb{Z}_{p}\}$ (55)
for $\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}$ and we define $\mathbf{X}^{\bm{\gamma},\bm{\zeta}}=\Psi(\mathcal{K}^{\bm{\gamma},\bm{\zeta}})$. Considering all these assumptions, we now propose the construction of the 2-D ZCAC and 2-D ZCACS.
Theorem 1.
Proof.
@@ 3.2 Main Construction
[…]
@@
$\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})=2^{m+1}p\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}+1}.\end{split}$ (57)
@@
$\begin{split}C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})=&\sum_{i=0}^{L_{1}-\tau_{1}-1}\sum_{j=0}^{L_{2}-\tau_{2}-1}\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}}_{i,j}{\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}_{i^{\prime},j^{\prime}}}^{*}\\ =&\sum_{i=0}^{L_{1}-\tau_{1}-1}\sum_{j=0}^{L_{2}-\tau_{2}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i^{\prime}}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j^{\prime}})},\end{split}$ (58)
where $i^{\prime}=i+\tau_{1}$, $j^{\prime}=j+\tau_{2}$. Let $1+\sum_{\alpha=1}^{k}m_{\alpha}=\epsilon$ and $m+1=\vartheta$. We use the following notations:
  1. 1.
@@
$\begin{split}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j}^{\prime})}=&\omega_{q}^{(1-j_{\epsilon})\left\{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})+a_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+a_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})\right\}}\\ &\times\omega_{q}^{j_{\epsilon}\left\{a_{\mathbf{s}_{2}^{1},\mathbf{t}_{2}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})+b_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+b_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})\right\}}\\ &\times\omega_{q}^{-(1-j^{\prime}_{\epsilon})\left\{a_{\mathbf{s}_{1}^{2},\mathbf{t}_{1}^{2}}^{\bm{\gamma}}(\vec{j}^{\prime}_{\xi})+a_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+a_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right\}}\\ &\times\omega_{q}^{-j^{\prime}_{\epsilon}\left\{a_{\mathbf{s}_{2}^{2},\mathbf{t}_{2}^{2}}^{\bm{\gamma}}(\vec{j}^{\prime}_{\xi})+b_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+b_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right\}},\end{split}$ (59)